# EB-MBD: Emerging-Barrier Model-Based Diffusion for Safe Trajectory Optimization in Highly Constrained Environments

### v1

@@ Abstract
We propose enforcing constraints on Model-Based Diffusion by introducing emerging barrier functions inspired by interior point methods. We show that constraints on Model-Based Diffusion can lead to catastrophic performance degradation, even on simple 2D systems due to sample inefficiency in the Monte Carlo approximation of the score function. We introduce Emerging-Barrier Model-Based Diffusion (EB-MBD) which uses progressively introduced barrier constraints to avoid these problems, significantly improving solution quality, without the need for computationally expensive operations such as projections. We analyze the sampling liveliness of samples each iteration to inform barrier parameter scheduling choice. We demonstrate results for 2D collision avoidance and a 3D underwater manipulator system and show that our method achieves lower cost solutions than Model-Based Diffusion, and requires orders of magnitude less computation time than projection based methods.
@@ I Introduction
Motion planning problems often have constraints such as collision avoidance or dynamic safety, which gradient-based methods have mature ways to incorporate. In the MBD diffusion framework, infeasible solutions result in zero-probability regions on the support of the distribution. Although various methods for enforcing constraints on learning-based diffusion have been proposed[9, 10, 11, 12], they do not translate to MBD due to its reliance on Monte Carlo sampling which is also affected by the constraints. When infeasible solutions cover a large part of the solution space, the score estimate’s value is dominated by rare events on the Monte Carlo proposal distribution, degrading performance.
Then contributions of this paper are:
@@ II-A Diffusion Models
While various attempts have been made to run diffusion models with constraints, they usually involve modifying the training process [9, 10]. [10] uses log barriers for constrained diffusion to modify the diffusion step size through the Riemannian metric induced by the barrier’s Hessian. However, this method is limited to convex constraints. Alternatively, they may try to enforce constraints on a pre-trained model during the denoising steps [12, 13, 11]. Many of these approaches only focus on learned diffusion models and do not translate to Monte Carlo score estimation which bring additional challenges due to how they are affected by constraints. Additionally, many use computationally expensive operations like projections onto non-convex constraint sets [11, 12] or optimization problems [13] which are passed to numerical NLP solvers whose solution quality and computation time can depend heavily on the initial guess and cost functions.
@@ III-A Diffusion Models
$x_{s+1}=\sqrt{1-\beta}\,x_{s}+\beta_{s}z_{s},$ (1)
@@ III-B Model-based Diffusion
$p(x)\propto\exp\left(-\frac{1}{\lambda}J(x)\right)$ (4)
@@
However, this Monte Carlo approach poses problems in a constrained environment where a large part of the solution space violates constraints. We demonstrate that its performance catastrophically degrades as constraints become restrictive even on simple 2D systems. If the problem is heavily constrained, the target density $p(x)=0$ in much of the solution space and samples $\hat{x}_{i}$ are likely to have no contribution, leading to “dead” samples with no information. Figure 2 shows how increasing obstacle radius in an existing obstacle avoidance problem, which MBD otherwise performs well on.
@@ IV-C Emerging Barriers
Input: Noise schedule $\beta_{s}$, Barrier schedule $(\mu_{s},c_{s})$, Diffusion steps $S$, Cost function $f(x)$, Constraint $g(x)$, Temperature $\lambda$
@@
 $x_{s-1}\leftarrow\frac{1}{\sqrt{\alpha}}\left[x_{s}+(1-\bar{\alpha})\gamma\right]$
@@ IV-D Barrier analysis
Theorem 1.
Suppose assumptions 1-4 hold, then the probability that a sample drawn from $X_{s}$ is alive is lower bounded by
$\mathbb{P}(X_{s}\text{ alive})\geq\Phi\left(\frac{1}{\sigma_{s}}\left(\frac{\mu_{s+1}}{L_{J}}-\frac{1-\sqrt{\bar{\alpha}_{s-1}}}{\sqrt{\bar{\alpha}_{s-1}}}R+c_{s}-c_{s+1}\right)\right).$
Proof.
@@ V Experimental Results
We implement Emerging Barrier MBD using the JAX Python package and demonstrate results for both a 2D obstacle avoidance environment, and a 3D high DOF underwater mobile manipulator system using the MuJoCo MJX as the underlying simulator [^1] Code can be found at https://github.com/LaVieEstDure/emerging_mbd. MJX allows simulation rollouts to be parallelized on a GPU. All experiments were conducted on a PC with an i7-13700K, and RTX 3070 GPU and 32 GB of memory.
@@
We show the performance of EB-MBD on a simple 2D obstacle avoidance problem. The dynamics and cost are given by
@@
where $\xi_{r}$ is a target position, and $\texttt{sigmoid}(x)=\frac{1}{1+e^{-x}}$. $g(\tau)$ is the signed distance to the closest obstacle encountered over the trajectory. MBD struggles due to amount of dead samples.
@@ V-B Comparison to projection-based methods
Figure 6: Typical trajectories from MBD, EB-MBD (Ours), Projected MBD[12] and DPCC-MBD[11]. EB-MBD performs substantially better than projection methods and is the only one that reaches the proximity of the goal
@@
EB-MBD is able to produce significantly lower cost trajectories and lower terminal distance to the target end point, all while taking takes orders of magnitude less time. Unlike the projection methods, EB-MBD also does not require taking derivatives of rollouts or the constraint function.
@@
TABLE I: Comparison of EB-MBD with variations of MBD on the 2D obstacle avoidance problem over 50 trajectories
@@ V-C Underwater Vehicle Manipulator System
and $J(\tau)=\ell_{T}(\xi_{T+1})+\sum_{t=0}^{T}\ell(\xi_{t},u_{t})$, where $\xi_{w}$ is the wrist position component of $\xi$, $r_{t}$ is the target wrist position, $\Lambda$ is a diagonal weighting matrix and $q$ is an orientation quaternion. This problem is challenging due to its high dimensionality and the complex motion required; the base is required to move and multiple joints need to coordinate together without the inertia of the robot causing a collision later in the trajectory. Figure 7 (a) shows how EB-MBD is able to successfully plan for the motion and 7 (b) shows how MBD trajectories get stuck in a poor local minima when EB-MBD finds early trajectories through the obstacle which get pushed out to better trajectories over iterations. Table II shows how EB-MBD achieves lower mean cost and higher success rates – which we define as the percentage of time the end effector was inside the box obstacle. We were unable to run projection methods in a reasonable amount of computational time on the UVMS due to increase in complexity. We emphasize the scalability of EB-MBD as computational time only grows with the extra evaluations of $g(x)$, as opposed to the complexity of the non-linear program which often grows more rapidly for complex systems.
Figure 7: (a) With EB-MBD, the UVMS reaches the target point while avoiding the obstacle. (b) Intermediate end effector trajectories over a single inference of EB-MBD and MBD visualized, with curves increasing in linewidth and redness towards the end of the process.
@@
TABLE II: Comparison of EB-MBD with MBD on 3D UVMS system over 50 trajectories
@@ VI Conclusion
While EB-MBD exhibited strong performance in our experiments, it does have some limitations. EB-MBD does not guarantee good solutions and good performance relies on barrier schedules well tuned for the problem. If a poor schedule is chosen, the output solutions are can be entirely infeasible. Our analysis also assumes timescale separation of the diffusion process local convergence and the barrier emergence. Future work may involve adaptive barrier schedules that prevent permanently dead samples without additional barrier tuning, and further analysis of EB-MBD under milder assumptions.
@@
where we make make use of Assumption 1. Using Lipschitz continuity of $g(x)$ and using Appendix -A to write in terms of one dimensional Gaussian cumulative distribution function, we find
@@
$\qquad\qquad\qquad\left.\nabla g\left(\frac{x_{s}}{\sqrt{\bar{\alpha}_{s-1}}}\right)^{T}\varepsilon+c_{s}\geq 0\middle|x_{s})\right)$

### v2

@@ Abstract
We propose enforcing constraints on Model-Based Diffusion by introducing emerging barrier functions inspired by interior point methods. We demonstrate that the standard Model-Based Diffusion algorithm can lead to catastrophic performance degradation in highly constrained environments, even on simple 2D systems due to sample inefficiency in the Monte Carlo approximation of the score function. We introduce Emerging-Barrier Model-Based Diffusion (EB-MBD) which uses progressively introduced barrier constraints to avoid these problems, significantly improving solution quality, without expensive projection operations such as projections. We analyze the sampling liveliness of samples at each iteration to inform barrier parameter scheduling choice. We demonstrate results for 2D collision avoidance and a 3D underwater manipulator system and show that our method achieves lower cost solutions than Model-Based Diffusion, and requires orders of magnitude less computation time than projection based methods.
@@ I Introduction
Motion planning problems often have constraints such as collision avoidance or dynamic safety, which gradient-based methods have mature ways to incorporate. In the MBD framework, infeasible solutions result in zero-probability regions on the support of the distribution. Although various methods for enforcing constraints on learning-based diffusion have been proposed[8, 4, 13, 3], they do not translate to MBD due to its reliance on Monte Carlo sampling which is also affected by the constraints. When infeasible solutions cover a large part of the solution space, the score estimate’s value is dominated by rare events on the Monte Carlo proposal distribution, degrading performance.
The contributions of this paper are:
@@ II-A Diffusion Models
While various attempts have been made to run diffusion models with constraints, they usually involve modifying the training process [8, 4]. [4] uses log barriers for constrained diffusion to modify the diffusion step size through the Riemannian metric induced by the barrier’s Hessian. However, this method is limited to convex constraints. Alternatively, they may try to enforce constraints on a pre-trained model during the denoising steps [3, 21, 13]. Many of these approaches only focus on learned diffusion models and do not translate to Monte Carlo score estimation which brings additional challenges due to how they are affected by constraints. Additionally, many use computationally expensive operations like projections onto non-convex constraint sets [13, 3] or optimization problems [21] which are passed to numerical NLP solvers whose solution quality and computation time can depend heavily on the initial guess and cost functions.
@@ III-A Diffusion Models
$x_{s+1}=\sqrt{1-\beta_{s}}\,x_{s}+\beta_{s}z_{s},$ (1)
@@ III-B Model-based Diffusion
$p(x)\propto\exp\left(-\frac{1}{\lambda}J(x)\right),$ (4)
@@
However, this Monte Carlo approach poses problems in a constrained environment where a large part of the solution space violates constraints. We demonstrate that its performance catastrophically degrades as constraints become restrictive even on simple 2D systems. If the problem is heavily constrained, the target density $p(x)=0$ in much of the solution space and samples $\hat{x}_{i}$ are likely to have no contribution, leading to “dead” samples with no information. Figure 2 shows how increasing obstacle radius in an existing obstacle avoidance problem causes failure, which MBD otherwise performs well on.
@@ IV-C Emerging Barriers
Input: Noise schedule $\beta_{s}$, Barrier schedule $(\mu_{s},c_{s})$, Diffusion steps $S$, Cost function $J(x)$, Constraint $g(x)$, Temperature $\lambda$
@@
 $x_{s-1}\leftarrow\frac{1}{\sqrt{\alpha_{s}}}\left[x_{s}+(1-\bar{\alpha}_{s})\gamma\right]$
@@ IV-D Barrier analysis
Theorem 1
Suppose assumptions 1-4 hold, then $\mathbb{P}(X_{s}\text{ alive})$, the probability that a sample drawn from $X_{s}$ is alive, is lower bounded by
$\Phi\left(\frac{1}{\sigma_{s}}\left(\frac{\mu_{s+1}}{L_{J}}-\frac{1-\sqrt{\bar{\alpha}_{s-1}}}{\sqrt{\bar{\alpha}_{s-1}}}R+c_{s}-c_{s+1}\right)\right).$
Proof:
@@ V Experimental Results
We implement Emerging Barrier MBD using the JAX Python package and demonstrate results for both a 2D obstacle avoidance environment, and a 3D high DOF underwater mobile manipulator system using the MuJoCo MJX as the underlying simulator [^1] Code can be found at https://github.com/acfr/emerging_barrier_mbd. MJX allows simulation rollouts to be parallelized on a GPU. All experiments were conducted on a PC with an i7-13700K, and RTX 3070 GPU and 32 GB of memory.
@@
We show the performance of EB-MBD on a simple 2D obstacle avoidance problem. The dynamics and cost are
@@
where $\xi_{r}$ is a target position, and $\texttt{sigmoid}(x)=\frac{1}{1+e^{-x}}$. $g(\tau)$ is the signed distance to the closest obstacle encountered over the trajectory. MBD struggles due to dead samples.
@@ V-B Comparison to projection-based methods
Figure 6: Typical trajectories from MBD, EB-MBD (Ours), Projected MBD[3] and DPCC-MBD[13]. EB-MBD performs substantially better than projection methods and is the only one that reaches the goal
@@
EB-MBD is able to produce significantly lower cost trajectories and lower terminal distance to the target end point, all while taking orders of magnitude less time. Unlike the projection methods, EB-MBD also does not require taking derivatives of rollouts or the constraint function.
@@
TABLE I: Comparison of EB-MBD on the 2D obstacle avoidance over 50 trajectories
@@ V-C Underwater Vehicle Manipulator System
and $J(\tau)=\ell_{T}(\xi_{T+1})+\sum_{t=0}^{T}\ell(\xi_{t},u_{t})$, where $\xi_{w}$ is the wrist position component of $\xi$, $r_{t}$ is the target wrist position, $\Lambda$ is a diagonal weighting matrix and $q$ is an orientation quaternion. This problem is challenging due to its high dimensionality and the complex motion required; the base is required to move and multiple joints need to coordinate together without the inertia of the robot causing a collision later in the trajectory. Figure 7 (a) shows how EB-MBD is able to successfully plan for the motion and 7 (b) shows how MBD trajectories get stuck in a poor local minima when EB-MBD finds early trajectories through the obstacle which get pushed out to better trajectories over iterations. Table II shows how EB-MBD achieves lower mean cost and higher success rates – which we define as the percentage of time the end effector was inside the box without collision. We were unable to run projection methods in a reasonable amount of computational time on the UVMS due to increase in complexity. We emphasize the scalability of EB-MBD as computational time only grows with the extra evaluations of $g(x)$, as opposed to the complexity of the non-linear program which often grows more rapidly for complex systems.
Figure 7: (a) With EB-MBD, the UVMS reaches the target point while avoiding the obstacle. (b) Intermediate end effector trajectories over a single inference of EB-MBD and MBD, with curves increasing in linewidth and redness towards the end of the process.
@@
TABLE II: Comparison of EB-MBD on 3D UVMS system over 50 trajectories
@@ VI Conclusion
While EB-MBD exhibited strong performance in our experiments, it does have some limitations. EB-MBD does not guarantee good solutions and good performance relies on barrier schedules well tuned for the problem. If a poor schedule is chosen, the output solutions can be entirely infeasible. Our analysis also assumes timescale separation of the diffusion process local convergence and the barrier emergence. Future work may involve adaptive barrier schedules that prevent permanently dead samples without additional barrier tuning, and further analysis of EB-MBD under milder assumptions.
@@
where we make make use of Assumption 1. Using Lipschitz continuity of $g(x)$ and using Appendix -A to write in terms of one dimensional Gaussian cumulative distribution function,
@@
$\qquad\qquad\qquad\left.\nabla g\left(\frac{x_{s}}{\sqrt{\bar{\alpha}_{s-1}}}\right)^{T}\varepsilon+c_{s}\geq 0\middle|x_{s}\right)$